In a particular electroplating process, the metal being plated has a +2 charge and a molar mass of 108.117 g/mol. If 589.62 C of charge pass through the cell, how many grams of metal should be plated? Useful information: F = 96,500 C/mol e- Provide your response to two digits after the decimal.
Answer #1

Amount of material deposited=q/zF

Z=charge=2,F=96500 C mol-1

q=589.62C

Plated=589.62/(2×96500)=0.003 moles=0.003×108.117 g/mol=

0.33 g plated